OpenStudy (anonymous):

hello to all please help me figure out the range for this problem f(x)=sqrt x-3

OpenStudy (anonymous):

is it \(f(x)=\sqrt{x}-3\)?

OpenStudy (anonymous):

\[fx)=\sqrt{x-3}\]

OpenStudy (anonymous):

I believe that the domain is \[\ge 3\]